Express the definite integral as a limit of Riemann sums. DO NOT EVALUATE.

k (x² + 2 ) dx
Answer #1

Given: S (x + 4) om AKI 52 -3 , Xi = Q +isz = a +3, - Hence, р X Simit :) dx = lines forces and = lim ntoo Jiw